OpenStudy (anonymous):

convert -250 degrees to radians

OpenStudy (lgbasallote):

multiply by pi/180

OpenStudy (anonymous):

i got -5pi/9 and was wondering if that's the correct answer

OpenStudy (lgbasallote):

seems right to me...

OpenStudy (lgbasallote):

oh wait...

OpenStudy (lgbasallote):

\[\huge -250 \times \frac{\pi}{180} \implies 25 \cancel 0 \times \frac{\pi}{18\cancel 0}\] right?

OpenStudy (lgbasallote):

uhh there should be a negative sign there

OpenStudy (lgbasallote):

\[\huge -250^o \implies -\frac{25\pi}{18}\]

OpenStudy (anonymous):

Yeah, that looks good. I think i was just trying to over reduce

OpenStudy (anonymous):

DAMMIT PI! http://tauday.com/